An oscillating pendulum bulb has a mass 0.1kg. It's amplitude is 0.01 meter and it's period is 2sec. The constant w (omega) in the acceleration formula is?
1
Expert's answer
2019-08-14T13:23:04-0400

The angular frequency for a simple pendulum is given by:



"w=\frac{2\pi}{T}"


Where

  • Period T=2sec


Numerically evaluating

"w=\frac{2\pi}{2}--->w=\pi \frac{rad}{s}"


Then the angular frequency "w=\pi \frac{rad}{s}"


The equation of maximum acceleration is

"a_{max}=A*w^{2}"

Where:

Amplitude of motion A=0.01m

angular frequency of movement "w=\\pi \\frac{rad}{s}\n\n\u200b"